Inessential Projects

The Student Information Processing Board January 21, 1994

1 Introduction

So, now you’re a SIPB prospective and you want to be useful. Often, prospectives feel they don’t know enough to be useful on a project, and this couldn’t be further from the truth. There are lots of things someone with next to no knowlegde can do and leave the project having been productive and having learned a lot. Above all, do your best not to be discouraged.

2 Some General Guidelines

Getting involved in SIPB projects is not always the easiest task, and this document aims to remedy some of that. Specifically, here are some guidelines for how to get involved.

• Ask dumb questions. If you have questions, ask them. Figure out who is best at answering questions on particular topics. Some particular tips on asking questions:

– Ask in person. Complicated explanation of things over zephyr makes things more confusing. Learning in person is often much more effective.

– Start at the bottom. If you are trying to figure how something works, start with the very beginning. It can be very difficult to learn something if it turns out one initial assumption was wrong.

• Ask members what they are working on. Most members are happy to explain what particular things they do in SIPB. Sometimes they may be busy, but eventually they are prone to be willing to tell you about it. This helps you figure out what kinds of things you are specifically interested in. If you are interested, ask them what you can read to learn more.

• Volunteer to do people’s work for them. If your being involved in somebody’s project will save them work, they will often be happy to help you learn more about it.

• Don’t be discouraged.

So, on to some more specific things.

3 Learning Stuff

Often prospectives don’t involve themselves in project because they personally don’t feel they know enough, and no matter how many times members say they do, they won’t believe it. So, here are some guidelines on teaching yourself some stuff so that your confidence level goes up. Remember though, often the best way to learn about something is just to dive in.

• Read documentation. Read SIPB and Athena documentation, and when you don’t understand something, ask someone about it. Come into the office, sit in the back and read and ask someone when you find something you are unsure of. When you think you understand, ask if your understanding is correct.

• Play around. Mess around on Athena. Ask people how they do things.

• Read net stuff. Read documents like ”Zen and the Art of Internet”, or ”The Big Dummy’s Guide to the Internet” or any of the dozens of cool documents out there. Ask a member to show you how to get at them.

• Work on a project. This is still one of the best ways to learn things. Ask questions.

Again, there is no minimum knowledge requirement or anything like it.

4 Projects

Ok, specific kinds of projects. Projects are changing all the time but there are some broad categories that remain the same, and often more appropriate for prospectives. Certainly, some kinds of projects require specific kinds of knowlege, but many require knowledge that you can easily learn on your own.

4.1 Documentation

One of the services the SIPB provides is documentation. It is often an ideal prospective project to work with SIPB documentation. To work on an existing SIPB document, one doesn’t need to be an expert on the topic, in fact, it is often useful not to be, for it allows one to notice important omissions in the document. Ask a member if you are interested in working on any specific documentation and start asking questions about what the document covers. Specific requirements and suggestions for a project of this sort are:

• Read the Document. Reading the document in question and possibly even taking notes on unclear parts is a definite necessity.

• Learn some LaTEX. All SIPB documentation is written in LaTEX, a text processing program, which you can obtain a working knowledge of by reading the SIPB document “Inessential LaTEX”.

• Ask questions. Ask lots of questions about the topic of the document so you can understand it well and write about it.

If you are interested in writing a new SIPB document, ask around to see if there is a need and ask members about what is necassary to begin doing so.

4.2 Ongoing Services

The SIPB has a number of ongoing services which have maintainers who may or may not have use for more maintainers.

4.2.1 News

Netnews at MIT is maintained by the SIPB. News maintenance is, unfortunately, not easy, but it is learnable. If you are interested in helping with news maintenance, ask around the office to find out who the current news maintainers are, and speak with them. A small warning though, news does actually require a fair amount of knowledge of UNIX, but if you are interested, ask anyway.

4.2.2 WorldWide Web

The SIPB maintains a WorldWide Web server. This is a distributed information system accessible via programs like Mosaic and lynx. Play around with the Web, learn HTML (the hypertext markup language) and if you are interested, ask around and speak to the current maintainers of the web server.

4.2.3 RTFM services

The SIPB maintains a programed named RTFM.mit.edu which contains ftp archives of FAQ (frequently asked questions) files, an addresses database and numerous other interesting things. If you are interested in maintaining services on this machine talk to people in the office and people can point you to the individuals who currently oversee that.

4.2.4 Random Other Services

There are various other services the SIPB supports, such as the Webster dictionary server, discuss, AFS, and many other things I’m surely not thinking of. Ask around.

4.2.5 New Services

If you think you are interested in maintaining a new service of some sort. Talk to people about the feasability of it. See if anyone else is interested in helping and pursue it.

4.3 Software

The SIPB maintains a lot of software. Look in the SIPB locker. If you use a program in the SIPB locker and think you may be interested in doing work with a that particular piece of software, ask around at meetings in the office and find out what would be involved.

5 What makes it a SIPB project though?

Here’s an analogy that might help. It’s along the same lines as the difference between doing an assignment in your high school physics class, and doing random physics problems in your physics textbook because you think they’re fun. Your teacher isn’t likely to count the random physics problems as part of your physics assignment, even if they were harder, or you found them interesting, or whatever. Nor will they let the random problems replace the ones you were asked to do.

Now, SIPB projects are not an assignment, of course. But the other difference is that a project you do on your own is something you thought was necessary, and you’re responsible for. If you wrote a program, you’d keep it in your own homedirectory, etc. You wouldn’t need anyone to think it was a good idea besides you. A SIPB project, on the other hand, is often something related to a service SIPB already provides, or wants to provide. It’s approved or suggested by SIPB in some way.

Hence, it’s a SIPB project to offer to put information about the IAP classes SIPB is sponsoring into xmosaic, for example. It would also be a SIPB service to say to some SIPB members, ”hey, there’s this program that i’ve seen elsewhere that i have the source to that isn’t on athena, and i think it would be cool if SIPB supported it. I’d like to compile it and maintain it.” Then you could offer to do it at a meeting, and if people though it was cool for SIPB to support it, you could work on it as a SIPB project.

6 What am I supposed to do in the office? Study Manuals?

Do you need to know C? If you learn C what do you do with it?

Sitting in the office and studying manuals is certainly not the only thing you can do. Especially if you’re like me, and don’t learn much from staring at a manual, or necessarily find it very interesting.

Knowing C is also not necessary. It’s a useful programming language, but SIPB is not ”Student groupof-people-who-like-to-read-manuals board” or ”Student C-programmers-incorporated Board.” There are a number of things SIPB tries to do, which are our purpose for existing. Basically, SIPB members want to improve computing at MIT by helping users, providing them with services that might not be available to them otherwise, and providing useful documentation on computing. (i’m missing lots of other specific examples...) For example, SIPB gave out accounts on old crufty systems back before athena existed. SIPB offered a machine you could dial up into from a modem before athena had dialup servers.

Prospectives aren’t expected to be experts on anything when they become prospectives. (And members aren’t expected to be experts on everything when they become members...) Nor are they expected to do all their learning in the SIPB office. If you want to find out how to program in C, you certainly don’t have to pick up SIPB’s copy of K&R, sit at a SIPB machine and read it, and write code. If, however, you wanted to learn about running linux on a PC and doing athena things on it, it might be useful to try asking one of the SIPB members who built most of athena for quiche-lorraine, the PC in the SIPB office which runs linux. The office is a resource, not an obligation.

There’s no rule that says you have to be working hard at learning something or writing some program whenever you come into the office. A lot of prospectives come into the office as a place to log in instead of in a cluster, to read mail, send zephyrs, etc. It’s generally a good thing to sit in the office and try to answer questions when people ask them, too, but you don’t have to answer them, or know the answers. If you’re totally unwilling to ever try to help a user who comes into the office, even when you do know how to help them, people probably would think this is bad, though.

7 I’m still confused whenever I’m in the office

Being a prospective means that you are interested in what goes on in the office, and that it IS your place to ask things, either about a conversation that’s going on, or at random. If you do want to know what people are talking about, it is perfectly reasonable to ask. If it’s none of your business, someone will tell you that. Or if they explain and you’re totally lost, you can say you’re lost. If they were talking about something important that they need to keep talking about, then they’ll say ”ask me later.” if they have time to explain it then, great. You’ll learn something. Listening to a conversation is a contribution, even if you know nothing about it, or aren’t an expert...because YOU learn.

As for people who seem to be working on private, secret projects– If i’m working on something at my machine, typing away, and you ask me what i’m doing, I would say something similar. If, for example, I were writing top secret code for the CIA and i’d be shot if anyone knew about it (which i don’t think is the case for anyone around here), I would tell you to go away. If I was sending mail to my best friend at another college, I would say that, and you probably wouldn’t be interested. If I was working on something for a SIPB project, or something you would find interesting, I would probably say that. Or if I were working on it, and I was in the middle of getting something done and wanted to finish my train of thought, I might say that, too, and ask you to wait. None of these things are bad...and a lot of them might result in you finding something out. But you don’t have to do anything liek that. If you don’t want to learn by asking people about random things, you don’t have to. You can decide what you do want to learn, and work on learning it.

8 So, members have to dedicate their lives to SIPB?

Of course people aren’t expected to dedicate their lives to SIPB. They aren’t even expected to dedicate n hours a week to it. It’s your own choice. If you don’t seem to want to devote anything at all to SIPB, it’s likely that you probably won’t become a member anyway. And if you don’t have time to devote for a while, you might become inactive if you are already a member, or wait a while to be considered for membership if you’re a prospective. You’d still be a member or a prospective, you just would not be considered someone who’s currently doing much with SIPB.

Frequently prospectives will spend too much time in the office, at the expense of performance in class. There are a decent number of people in SIPB who think that doing things for SIPB and other random computer things are more fun than their classes, and let this take a toll on their classes. This is a bad thing, though. It’s not a requirement for membership in SIPB, and most SIPB members try to encourage each otehr to go home and tool when they need to.

9 Is this any fun?

If you don’t enjoy helping people or learning about things, then there’s not much SIPB can do about it. Don’t try to be a prospective if you don’t enjoy what you’re doing, maybe you’re making the wrong decision.